OpenStudy (hannahwaller5):

Will Give Medal!!! An expression is shown below: f(x) = –16x2 + 24x + 16 Part A: What are the x-intercepts of the graph of f(x)? Show your work. Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph.

OpenStudy (anthonyym):

x-intercepts are the zeros of the function, where the graph crosses the x-axis, aka at y=0. Therefore set the equation equal to zero. f(x) is another way of saying y. \[y=–16x^2+24x+16\] \[0 = –16x^2 + 24x + 16\]

OpenStudy (hannahwaller5):

Yes.. that i know how to do... I know also that you then have to find the GCF and then factor that out... when you do this, i get 8(-2x^2 +3x + 2) (i think this is right...?) Then i know you are supposed to multiply -2 and 2 (you get 4) then find the factors of 4 and see which of the two factors you find that you can add together in order to find the middle term (3) I THINK this is how you do it... am i right? If i am right i have no idea what i am doing wrong... if i am wrong i have no idea what to do...

OpenStudy (hannahwaller5):

Oh ok... never saw how to do that before lol but it makes since... what do i do after that?

OpenStudy (anthonyym):

Actually you can divide the whole equation by 8 since it's equal to 0. Divide both sides by 8. (0 divided by 8 is still 0). It becomes: 0 = -2x^2 + 3x + 2 (sorry divided wrong previously)

OpenStudy (hannahwaller5):

Oh ok thats what i got when i did the GCF.. so now i am confused on what to do next...

OpenStudy (anthonyym):

Now you can factor. You need to numbers that add to the coefficient of x (which is 3) and multiply to the product of the constant and coefficient of x^2 (product is -4).

OpenStudy (hannahwaller5):

ok... sorry im a little confused

OpenStudy (anthonyym):

This is how I like to organize it. 

OpenStudy (anthonyym):

Do you understand how I got the 3 and 4?

OpenStudy (hannahwaller5):

I think that 3 is the middle term and you got 4 by multiply 2 times 2...?

OpenStudy (anthonyym):

Yes. Now the two numbers you found, 4 and -1, you put into the equation like this: 0 = -2x^2 + 3x + 2 0 = -2x^2 + 4x -x + 2 Notice the 3x is broken down into 4x and -1x, that's where the numbers come in; they're the coefficients now.

OpenStudy (hannahwaller5):

Oh ok... but why would you put 4 and 1 in place of 3?

OpenStudy (hannahwaller5):

Nevermind i see what you wrote... ok so now we solve the problem regularly?

OpenStudy (hannahwaller5):

If i did i would get -2x(x+2) -1(x+2)... am i right?

OpenStudy (anthonyym):

Not completely..

OpenStudy (hannahwaller5):

ok...what did i do wrong

OpenStudy (anthonyym):

You have 0 = -2x^2 + 4x -x + 2 2x(x+2) -1(x+2) If you distribute back 2x * x does not equal the first term -2x^2. And -1 * 2 does not get 2

OpenStudy (hannahwaller5):

Oh i see... so what did we do wrong because thats how you would solve it... i think

OpenStudy (anthonyym):

It should have some negatives in there... -2x^2 + 4x - x + 2 You chose to take out 2x. 2x(-x+2) - x + 2 And a rule of factoring further is for the second part to have (-x+2). The only way is to factor out a positive 1. 2x(-x+2) +1(-x+2)

OpenStudy (hannahwaller5):

Oh... ok i see i shouldnt have put the negatives on the outside of the parenthesis... So after that normally to find the x's (like it says to do in part a) i would then take 2x and 1 and put them together in a parenthesis and then thake the (-x+2) So i would have (2x+1)(-x+2) But now i dont understand what to do next...

OpenStudy (anthonyym):

Yes and that equals 0. I forgot to include the 0 in the others previously. 0 = (2x+1)(-x+2) So what makes the right side equal 0? Since it's multiplication, either one of them has to equal 0. So, 0 = 2x + 1 and 0 = -x + 2

OpenStudy (hannahwaller5):

im confused...

OpenStudy (anthonyym):

my bad it should be or instead of and 0 = 2x + 1 or 0 = -x + 2

OpenStudy (anthonyym):

hm which part?

OpenStudy (hannahwaller5):

the part where you said So what makes the right side equal 0? Since it's multiplication, either one of them has to equal 0.

OpenStudy (anthonyym):

That's the zero product of multiplication. 0 times anything is 0. So since (2x+1)(-x+2) has to equal 0, either 2x+1 has to equal 0 or -x + 2 has to equal 0. So in the end you get two solutions (two zeros) because of the "or" part

OpenStudy (hannahwaller5):

Oh yea thats right ok... so to make (2x+1) equal zero we would have to do 2x + 1 - 3 = 0 - 3 then the x intercept would be (-3,0) I think...

OpenStudy (anthonyym):

You're just solving the equation 2x + 1 = 0 Subtract 1 from both sides, then divide by 2

OpenStudy (hannahwaller5):

If i did that i would get x = -0.5...

OpenStudy (anthonyym):

Yes that's correct. That's one of the solutions. You also have -x+2 = 0

OpenStudy (hannahwaller5):

2

OpenStudy (hannahwaller5):

x=2

OpenStudy (anthonyym):

Correct

OpenStudy (hannahwaller5):

YAY!!!!

OpenStudy (hannahwaller5):

What about this... if you have time Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work.

OpenStudy (anthonyym):

Since there's a negative in front (-16) the graph will slope downward like an upside down v ( like a tepee). Therefore it will have a maximum since there's the tip at the top and the bottom (range) goes down to infinity and won't end.

OpenStudy (anthonyym):